OpenStudy (anonymous):

Which statement is true for the expression 6a + 15b − 9c?

OpenStudy (anonymous):

It can be written as 2(3a + 7b − 4c) by using associative property of multiplication. It can be written as 3(2a + 5b − 3c) by using associative property of addition. It can be written as 2(3a + 7b − 4c) by using distributive property of multiplication. It can be written as 3(2a + 5b − 3c) by using distributive property of addition.

OpenStudy (shane_b):

If you factor a 3 out of each term what expression do you end up with?

OpenStudy (anonymous):

umm the last one?

OpenStudy (shane_b):

No :) To factor a 3 out of each term it would look like this: \[\frac{6a}{3}+\frac{15b}{3}-\frac{9c}{3}=2a+5b-3c\]and since you divided everything by 3, you have to balance things out by multiplying this entire expression by 3. That leaves you with: \[3(2a+5b-3c)\]

OpenStudy (anonymous):

it is the last one D rite?

OpenStudy (shane_b):

Sorry, yes it's d.

OpenStudy (anonymous):

lol thanks !

OpenStudy (shane_b):

Good luck :)